OpenStudy (anonymous):

*Advance Algebra With Financial Applications* Keith owes $4,512 on a credit card with a 17.3% interest rate compounded monthly. What is the monthly payment he should make in order to pay off this debt in 12 months, assuming he does not charge any more purchases with the card? $915.50 $74.54 $376.00 $412.16 ------------ I believe the answer is A. Am I correct?

OpenStudy (anonymous):

m=month x= montly payment 4512-x=current_m current_m+current_m*17.3=current_(m+1) current_12=0 Something like that lol i would only know how to do it via computer.

OpenStudy (anonymous):

Hmmm

OpenStudy (anonymous):

i can make your bead rock

OpenStudy (anonymous):

you can do like the first few months and see if you notice a pattern

OpenStudy (anonymous):

I'm still confused haha

OpenStudy (anonymous):

o(m)=o(m-1)+x+i(o(m-1)-x)

OpenStudy (anonymous):

i think payment per month and interest are constants lets call it. x and i. And also initial debt is constant and i wall call it D month(m) 1 2 3 4 total paid(tp) x x*2 formula is x*m int owed(o) (ID-x)+i*(ID-x) o(1)-x+i(o(1)-x) o(2)-x+i(o(2)-x) 1 2 3 o(3)-x+i*(o(3)-x) 4 seems formula is o(m)=o(m-1)+x+i(o(m-1)-x)